OpenStudy (anonymous):

solve the integral x/(3x+5)

OpenStudy (mr.math):

Use a substitution \(u=3x+5\), this is one way.

OpenStudy (anonymous):

wont work !!!

OpenStudy (anonymous):

you will still have x's left ...

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

what do you mean solve for x ?

OpenStudy (mr.math):

\[u=3x+5 \implies x=\frac{1}{3}(u-5) \implies dx=\frac{1}{3}du.\]

OpenStudy (anonymous):

mr math is writing it now i am sure, but solve \[u=3x+5\] for x

OpenStudy (anonymous):

and then ? i tried polynomil division but couldn't do it ...

OpenStudy (mr.math):

Using the substitution, your integration becomes: \[\frac{1}{9}\int \frac{u-5}{u}du.\] But \(\frac{u-5}{u}=1-\frac{5}{u}\). You should be able to take it from here.

OpenStudy (mr.math):

Is that good enough?

OpenStudy (anonymous):

thank you !!! =]

OpenStudy (mr.math):

Glad to help :-)